use the following equation, where x and y are the measured in meters, to model a cross section of the Barringer Meteor crater, near Winslow, Arizona
y=1/1800(x-600)(x+600)
1) assuming the lip of the crater is at y=0, how wide is the crater?
2) what is the depth of the crater?

Answer by stanbon(75887)

use the following equation, where x and y are the measured in meters, to model a cross section of the Barringer Meteor crater, near Winslow, Arizona
--------------------------
y=1/1800(x-600)(x+600)
1) assuming the lip of the crater is at y=0, how wide is the crater?
(1/1800)(x-600)(x+600) = 0
(x-600)(x+600 = 0
Positive solution:
x = 600 meters
======================
2) what is the depth of the crater?
Assuming the crater is deepest in the center, let x = 300 meters;solve for y.
y = (1/1800)(300-600)(300+600)
y = (1/(1800)*-300*900
y = -150 meters
============================
Cheers,
Stan H.